OpenStudy (anonymous):

Samantha loves to download new music. She originally had 52 songs but plans to purchase 2 new songs each week. She wants to know how many songs she will have after 93 weeks. Which equation should she use? a93 = 93 + 2(52 − 1) a93 = 93 + 2(52 + 1) a93 = 52 + 2(93 + 1) a93 = 52 + 2(93 − 1)

OpenStudy (yanasidlinskiy):

For this question, I would recommend people like @whpalmer4 @tkhunny and @thomaster

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (tkhunny):

a93 = 93 + 2(52 − 1) a93 = 93 + 2(52 + 1) a93 = 52 + 2(93 + 1) a93 = 52 + 2(93 − 1) Originally 52 - Get rid of the first 2 a93 = 52 + 2(93 + 1) a93 = 52 + 2(93 − 1) 2 each week. - Annoying. That doesn't help. a93 = 52 + 2(93 + 1) a93 = 52 + 2(93 − 1) We're not buying any new ones in the first week, so purchases are one down. a93 = 52 + 2(93 − 1) I think we made it.

OpenStudy (anonymous):

Thank you soooo much!!!

OpenStudy (tkhunny):

Just think your way through it - one piece of information at a time. If you run out of information, think harder!!

OpenStudy (anonymous):

Every day, there are 5 times more likes on an internet video of a dog which is modelled by the function c(n) = (5)n − 1, where n is the number of days since the video posted. On the first day, there were 103 likes. What is the function that shows the number of likes each day? c(n) = (5)(103)(n − 1) c(n) = (103)n − 1 c(n) = 103(5)n − 1 c(n) = (5)103 − 1

OpenStudy (anonymous):

For this one I think its either A or D.

OpenStudy (tkhunny):

The notation is NOT helpful. You must do something to indicate and exponent. c(n) = 5^n - 1 = \(5^{n}-1\) Then \(c(103) = 5^{103} - 1\) - But, that doesn't quite work. You might need \(c(n) = 103\cdot 5^{n-1}\). This gives \(c(1) = 103\cdot 5^{1-1} = 103\cdot 1 = 103\) and I think we have it.